graph the line.
x-4y=-8
here is what I am getting so wanting to check it
-4y=-8-x
y=-2/-1

Answer by wilft1(217)

when you graph the line, you want to create a column for x and y
in the equation, substitute x or y for a number and then solve for the other variable, example...

x - 4y = -8
substitute x for 2, then solve for y
2 - 4y = -8
x = 2 and y = 2.5, make sure to make them an ordered pair (2,2.5)
but to graph it, your going to need to do this 3 times, i say 3 times just so you know your math is right and the 3 points make a straight line
lets say that x = 4
4 - 4y = -8
x = 4 and y = 3 (4,3)
one last time, lets say that x = -2
-2 - 4y = -8
x = -2 and y = 1.5 (-2,1.5)
so just go ahead and graph the 3 points, you'll notice they make a straight line